OpenStudy (anonymous):

i really need help can u Solve -4(x + 10) - 6 = -3(x - 2)

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 @familyguy123

TheSmartOne (thesmartone):

Do you know the distributive formula?

TheSmartOne (thesmartone):

Use this to solve your question and show your work so I can see how far you can go and/or where you ma have made the mistake: \(\sf a(b+c)=ab+ac\) \(\sf a(b-c)=ab-ac\)

OpenStudy (luke69603):

its 52

OpenStudy (luke69603):

i mean negative 52